### v1

# The loop equations for noncommutative geometries on quivers

Abstract.
We define a path integral over Dirac operators that averages over noncommutative geometries on a fixed graph, as the title reveals, partially using quiver representations. We prove algebraic relations that are satisfied by the expectation value of the respective observables, computed in terms of integrals over unitary groups, with weights defined by the spectral action. These equations generalise the Makeenko-Migdal equations—the constraints of lattice gauge theory—from lattices to arbitrary graphs. As a perspective, these constraints are combined with positivity conditions (on a matrix of parametrised by composition of Wilson loops). A simple example of this combination known as ‘bootstrap’ is fully worked out.
1. Introduction
Before discussing our problem in its due context, we describe it aridly and postpone its motivation for Section 1.1. For integers $n,N>1$, fix a polynomial $S\in\mathbb{C}_{\langle 2n\rangle}=\mathbb{C}{\langle u_{1},u_{1}^{*},u_{2},u_{2}^{*},\ldots,u_{n},u_{n}^{*}\rangle}$ in noncommutative $u$-variables satisfying $u_{j}u_{j}^{*}=1=u_{j}^{*}u_{j}$ for $j=1,\ldots,n$. Consider a family of integrals of the type
$I_{\beta}=\int_{\mathrm{U}(N)^{n}}\operatorname{Tr}\beta(U_{1},U^{*}_{1},\ldots,U_{n},U_{n}^{*})\mathrm{e}^{N\operatorname{Tr}S(U_{1},U^{*}_{1},\ldots,U_{n},U_{n}^{*})}{\mathrm{d}}U_{1}{\mathrm{d}}U_{2}\cdots{\mathrm{d}}U_{n},\quad\beta\in\mathbb{C}_{\langle 2n\rangle},$ (1.1)
with each factor ${\mathrm{d}}U_{i}$ being the Haar measure on $\mathrm{U}(N)$. Assuming that $\operatorname{Tr}S$ is real-valued over the whole integration domain, we derive the loop equations, that is to say, algebraic relations among the integrals $\{I_{\beta}\}_{\beta\in\mathcal{I}}$ parametrised by a certain family $\mathcal{I}\subset\mathbb{C}_{\langle 2n\rangle}$. This type of integrals has been considered by physicists in the context of lattice gauge field theory. In mathematics, integrals over the unitary group are relevant in the context of Weingarten-calculus [Col03], developed mainly by Collins and collaborators (e.g. [CS06, CGL24]).
1.1. Motivation: Random matrix theory and noncommutative geometry
Our interest in integrals of the type (1.1) emerges from Connes’ noncommutative geometrical [Conn94] approach to fundamental interactions, in which geometric notions are mainly governed by a self-adjoint operator $D$ named after Dirac. In this setting, the physical action $S(D)$ is claimed to depend only on (the spectrum of) $D$ and is known as spectral action [CC97]. Our problem is to evaluate the moments that the spectral action yields via
$\mathbb{E}[h(D)]=\frac{1}{\mathcal{Z}}\int_{\text{Dirac}}h(D)\mathrm{e}^{-S(D)}{\mathrm{d}}D,\qquad\mathbb{E}[1]=1,\quad h(D)\in\mathbb{R},$ (1.2)
for an ensemble of Dirac operators $D$ (the normalisation condition defines $\mathcal{Z}$). Of course, this requires to have defined the measure ${\mathrm{d}}D$ over such ensemble of Dirac operators $D$. (The original problem, as it appears in [CM08, Sec. 19], considers also fermions, as it has been recently addressed in [KPV24], but which we do not include yet.)
Part of the relatively vivid interest in the problem (1.2) during the last decade is due to the reformulation [Bar15] of fuzzy spaces [^1] We do not aim at a comprehensive review here, for fuzzy spaces see e.g. [StSz08] and the works of Rieffel [Rie10, Rie10, Rie23] (and references therein) that address, from diverse mathematical angles, the rigorous convergence of matrix algebras to the sphere. We are also not reviewing all the quantisation approaches either; for a Batalin-Vilkovisky approach: cf [IvS17] for Tate-Koszul resolutions applied to a model of $2\times 2$-matrices and [GNS22, NSS21] for the homological-perturbative approach to Dirac-operator valued integrals. as finite-dimensional spectral triples. This led to the application of random matrix theory tools [AK24, Pér22a, KP21, Pér21, Pér22b, HKPV22] that followed to the first numerical results [BG16]. All these works deal with multimatrix interactions that include a product of traces (as opposed to the ordinary interactions that are single a traces of a noncommutative polynomial).
Independently, [vS11, Cor. 19] Taylor-expands the spectral action as a matrix model of the form $V(M)=\sum_{l=1}^{\infty}\sum_{i_{1},i_{2},\ldots,i_{l}}F_{i_{1},i_{2},\ldots,i_{l}}M_{i_{1},i_{2}}M_{i_{2},i_{3}}\cdots M_{i_{l},i_{1}}$. This expansion was shown in [vNvS21] to be convergent and, taking some elements of [CC06] and with own techniques, to possess a neat reorganization in terms of universal Chern-Simons forms and Yang-Mills forms integrated against $(B,b)$-cocycles that do depend on the geometry. The resulting model $V(M)$ above goes beyond the solved [GHW20, BHGW22] generalisations of the Kontsevich matrix model [Konts92] (in which only the unitary invariance of the quadratic term is broken) known as Grosse-Wulkenhaar model [GW14].
These two independent approaches portend a symbiosis between random matrix theory and noncommutative geometry. Both the multiple trace interactions and the unitary broken interactions could motivate (if they have not yet) new developments in random matrix theory. And vice versa, the path-integral quantisation (1.2) of noncommutative geometries seems hopeless without the intervention of random matrix theory [^2] The only alternative known to the author is the use of Choi-Effros operators systems [CvS21, CvS22] (cf. also [DLL22]) that emerge when one assumes (or rather, when one accepts) that only a finite part of the Dirac spectrum is measurable. The price to pay is nonassociativity. .
1.2. Ensembles of unitary matrices in noncommutative geometry
The interaction between these two disciplines has taken place in hermitian grounds. In this letter, integrals over Dirac operators boil down to ensembles of unitary matrices (they are also unitary-invariant, like ordinary hermitian matrix ensembles, but unitary ensembles integrate over unitary random matrices). These can be considered as an approach to average over ‘noncommutative geometries on a graph’. When the graph is provided with additional structure, it might be grasped as a discretisation of space. For instance, edges would carry a representation, and vertices equivariant maps, in a spin network approach; here, we refrain from including information that associated to gravitational degrees of freedom and address exclusively the problem of gauge interactions. The background geometry is therefore fixed and the finiteness of the unitary groups appearing is not a shortage of the theory; as a caveat, they are not to be interpreted as a truncation of infinite-dimensional symmetries (but to be compared with the unitary structure group of Yang-Mills, for example).
Representation theory does still play a role, but rather in the context of quiver representations in a certain category that emerges from noncommutativity geometry as exposed in [Pér24] after the fundamental ideas of [MvS14].
We can now restate the aim of this letter as follows:
Define a partition function for noncommutative geometries on a graph—that is, define a measure over all Dirac operators—and prove the algebraic relations that the respective observables shall satisfy. Such quantities have the form $I_{\beta}$ as in eq. (1.1) and are called Wilson loops (although not each $I_{\beta}$ is a Wilson loop).
(Proper definitions follow in the main text.) Such relations generalise the Makeenko-Migdal equations, the loop equations in lattice gauge theory. After introducing the setting in Section 2, we prove the main result in Section 3 and conclude with a fully worked-out application that mixes the loop equations with positivity conditions of a certain matrix (‘bootstrap’) in Section 4.
2. Quiver representations and noncommutative geometry
We call quiver $Q$ a directed multigraph. Since $Q$ is directed, there are maps $s,t:Q_{1}\rightrightarrows Q_{0}$ (from the edge-set $Q_{1}$ to the vertex-set $Q_{0}$) determining the vertex $s(e)$ at which an edge $e$ begins, and the one $t(e)$ where it ends. Multiple edges $e,e^{\prime}\in Q_{1}$ and self-loops $o_{v}\in Q_{0}$ at a certain vertex $v\in Q_{0}$ are allowed, namely $\{s(e),t(e)\}=\{s(e^{\prime}),t(e^{\prime})\}$ as sets, and $s(o)=t(o)=v$, respectively.
One interprets a quiver $Q$ as a category whose objects are $Q_{0}$. The morphisms $\hom_{Q}(v,w)$ are the paths from $v$ to $w$, namely edge-sequences $\gamma=(e_{1}e_{2}\cdots e_{n})$ with $e_{1},\ldots,e_{n}\in Q_{1}$ and $s(\gamma)=s(e_{1})=v$, and $t(\gamma)=t(e_{n})=w$ as well as $t(e_{j})=s(e_{j+1})$ for $j=1,\ldots,n-1$. We shall write $\gamma:v\to w$ if $v=s(\gamma)$ and $w=t(\gamma)$ and call $\ell(\gamma)=n$ the length of $\gamma$. The path $\gamma$ with reversed order is denoted by $\bar{\gamma}=(e_{n}e_{n-1}\cdots e_{2}e_{1})$ (not to be confused with the inverse morphism of $\gamma$). Obviously, unless otherwise stated, paths are directed, but it will prove useful to consider also paths in $\Gamma Q$, the underlying graph of the quiver ($Q$ with forgotten orientations). If $s(\gamma)=t(\gamma)$ we call $\gamma$ a loop. The space of loops [^3] We comment for sake of completeness, that the space of endomorphisms $\Omega_{v}(Q)=\operatorname{\mathrm{hol}}_{Q}(v,v)$ has as identity the constant zero-length path, which does play a role in the theory of path algebras while constructing an equivalence between the category of representation and modules of the path algebra [DW17], but here we do not need this explicitly. at $v$, is denoted here $\Omega_{v}(Q)$, that is $\Omega_{v}(Q)=\hom_{Q}(v,v)$, and $\Omega Q$ will denote the space $\cup_{v\in{Q_{0}}}\Omega_{v}(Q)$ of all loops.
A quiver exists essentially to be represented (otherwise one would say multidigraph) in a category $\mathcal{C}$. A $\mathcal{C}$-representation of $Q$ is by definition a functor from $Q$ to $\mathcal{C}$.
2.1. The spectral triple associated to a quiver representation
We restrict the discussion to finite dimensions and introduce the setting of [Pér24]. By definition, an object in the category ${\mathscr{pS}}$ of prespectral triples is a pair $(A,H)$ of a *-algebra $A$ faithfully *-represented, $\lambda:A\curvearrowright H$, in an inner product $\mathbb{C}$-vector space $H$ (*-represented means here, that $\lambda(a^{*})$ is the adjoint operator of $\lambda(a)$ for all $a\in A$). A morphisms in $\hom_{{\mathscr{pS}}}(A_{s},H_{s}\,;\,A_{t},H_{t})$ is a couple $(\phi,U)$ of an involutive algebra map $\phi:A_{s}\to A_{t}$ as well as a unitary map $U:H_{s}\to H_{t}$. As part of the definition, a morphism should in addition satisfy $U\lambda_{s}(a)U^{*}=\lambda_{t}[\phi(a)]$ for all $a\in A$.
In other words, a ${\mathscr{pS}}$-representation of $Q$ associates with each vertex $v$ of $Q$ a prespectral triple $(A_{v},H_{v})\in{\mathscr{pS}}$ and with any path $\gamma:v\to w$ a morphism $(\phi_{\gamma},\operatorname{\mathrm{hol}}\gamma):(A_{v},H_{v})\to(A_{w},H_{w})$ in such a way that if $\gamma=(e_{1}\cdots e_{2})$, then $\operatorname{\mathrm{hol}}\gamma=U_{e_{n}}\cdots U_{e_{1}}$ and $\phi_{\gamma}=\phi_{e_{n}}\circ\phi_{e_{n-1}}\circ\cdots\circ\phi_{e_{1}}$, where $\phi_{e_{j}}:A_{s(e_{j})}\to A_{t(e_{j})}$ and $U_{e_{j}}:H_{s(e_{j})}\to H_{t(e_{j})}$ form a ${\mathscr{pS}}$-morphism. We refer to $\operatorname{\mathrm{hol}}\gamma$ as the holonomy of $\gamma$.
If two vertices are connected by a path $\gamma$, notice that $\operatorname{\mathrm{hol}}\gamma$ is a unitarity and $\dim H_{s(\gamma)}=\dim H_{t(\gamma)}$. If $Q$ is connected there might be no (directed) path between two given vertices $v$ and $w$; it is however easy, if necessary by inverting some subpaths of a path $\tilde{\gamma}$ in $\Gamma Q$ that connects $v$ with $w$, to establish the constancy of the map $Q_{0}\ni v\mapsto\dim H_{v}:=N$; we call such constant $N=\dim R$, the dimension of the representation $R$, somehow abusively.
A spectral triple $(A,H,D)$ is a prespectral triple $(A,H)$ together with a self-adjoint element $D\in\operatorname{\mathrm{End}}(H)$, referred to as Dirac operator. (This terminology comes from the hard fact that, if certain operators are added to the [in that case, infinite-dimensional] spectral triple and if, together with $D$, such operators, satisfy a meticulous list of axioms, then $D$ is the spin geometry Dirac operator [Conn13]; see also [vS15] for an introduction geared to physicists).
Remark 2.1.
As a side note, it is possible to compute the space of all ${\mathscr{pS}}$-representations of $Q$. It was proven in [Pér24] that such space—which in fact forms the category of representations—can be described in terms of products of unitary groups subordinated to combinatorial devices called Bratteli networks. This will not be recalled here, since we will focus only on the spectral action. Nevertheless, it is important to observe that, in stark contrast with ordinary $\mathsf{Vect}_{\mathbb{C}}$-quiver representations, providing labels to the vertices is not enough to determine a ${\mathscr{pS}}$-quiver representation. The lifts of whole paths should exist, and this needs the compatibility of the maps $\phi_{v}$ at all vertices $v$, which in turn is what the so-called Bratteli networks guarantee (concretely *-algebra maps for $M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ for $m>n$ do not exist, and if a representation yields $A_{s(e)}=M_{m}(\mathbb{C})$ and $A_{t(e)}=M_{n}(\mathbb{C})$ for some edge $e$, a lift fails, cf. [Pér24, Ex. 3.16]). Despite this, we denote representations of quivers as $R=\{(A_{v},H_{v}),(\phi_{e},U_{e})\}_{v\in Q_{0},e\in Q_{1}}$ instead of $R=\{(A_{v},H_{v}),(\phi_{\gamma},\operatorname{\mathrm{hol}}\gamma)\}_{v\in Q_{0},\gamma\in\Omega Q}$, under the tacit assumption that lifts of whole paths exist.
We associate now a spectral triple to a given ${\mathscr{pS}}$-representation $R=\{(A_{v},H_{v}),(\phi_{e},U_{e})\}_{v\in Q_{0},e\in Q_{1}}$ of a connected quiver $Q$. We define the Dirac operator associated to $R$ as the matrix $D_{Q}(R)\in M_{\#Q_{0}}(\mathbb{C})\otimes M_{N}(\mathbb{C})$ with matrix entries $[D_{Q}(R)]_{v,w}\in M_{N}(\mathbb{C})$ in the second factor given by
$[D_{Q}(R)]_{v,w}=\bigg{(}\sum_{e\in s^{-1}(v)\cap t^{-1}(w)}U_{e}\bigg{)}+\bigg{(}\sum_{e\in t^{-1}(v)\cap s^{-1}(w)}U_{e}^{*}\bigg{)}\qquad(v,w\in Q_{0}).$ (2.1)
By construction, this operator is self-adjoint, and crucially for our purposes, the next collection
$\big{(}A_{Q}(R),H_{Q}(R),D_{Q}(R)\big{)}=\bigg{(}\bigoplus_{v\in Q_{0}}A_{v},\bigoplus_{v\in Q_{0}}H_{v},D_{Q}(R)\bigg{)}$ (2.2)
is a spectral triple.
2.2. The spectral action and the partition function
Given a polynomial $f(x)=f_{0}+f_{1}x^{1}+f_{2}x^{2}+\ldots+f_{d}x^{d}$ in real variables $f_{0},f_{1},\ldots,f_{d}\in\mathbb{R}$, and a quiver representation, the spectral action on a quiver reads $S(D)=\operatorname{Tr}_{H}f(D)$, where we abbreviate $D=D_{Q}(R)$. It is possible to compute the spectral action as a loop expansion in terms of generalised plaquettes $\gamma$ as follows
$\operatorname{Tr}_{H}f(D)=\sum_{k=1}^{d}f_{k}\,\sum_{v\in Q_{0}}\,\,\sum_{\begin{subarray}{c}\gamma\in\Omega_{v}(Q)\\ \ell(\gamma)=k\end{subarray}}\operatorname{Tr}\operatorname{\mathrm{hol}}\gamma,$ (2.3)
where $\operatorname{Tr}$ in the rhs is the trace of $M_{N}(\mathbb{C})$ with $\operatorname{Tr}1=\dim R=N$. The proof of eq. (2.3) is given in [Pér24], but the reader will recognise this formula as a noncommutative generalisation of the following well-known fact in graph theory: and $C_{G}$ denotes the adjacency matrix of a graph $G$, the number of length-$n$ paths in $G$ between two of its vertices, $i$ and $j$, is the entry $[C_{G}^{n}]_{i,j}$ of the matrix $(C_{G})^{n}$.
Now we turn to the space of ${\mathscr{pS}}$-representations of $Q$. Let $A_{v}=\oplus_{j=1}^{l_{v}}M_{n_{v,j}}(\mathbb{C})$ denote the algebra associated by $R$ to the vertex $v$ (so $l_{v}$ is the number of simple subalgebras of $A_{v}$). Let $r_{v,j}$ be the multiplicity of the action of the factor $M_{n_{v,j}}(\mathbb{C})\subset A_{v}$ on the Hilbert space $H_{v}$, that is $H_{v}=\oplus_{j=1}^{l_{v}}\mathbb{C}^{r_{v,j}}\otimes\mathbb{C}^{n_{v,j}}$ where $M_{n_{v,j}}(\mathbb{C})$ only acts non-trivially on $\mathbb{C}^{n_{v,j}}$ via the fundamental representation.
Once labels to the vertices are consistently444‘Consistently’ is quantitatively described in terms of certain transition matrices in [Pér24, Def. 3.13], and qualitatively means that the labels are such that involutive maps $A_{s(e)}\to A_{t(e)}$ always exist for each $e\in Q_{1}$, and several dimension-matching checks. The reader will note that we do not include the minimal amount of information in each group at the edges. The origin of the projective groups $\mathrm{PU}(n)$ is that $\mathrm{U}(n)$ acts via the adjoint action. assigned, the possible labels of an edge $e$ are parametrised by $\prod_{j=1,\ldots,l_{t(e)}}\mathrm{U}(n_{t(e),j})$ [Pér24, Lemma 3.5], and at each vertex it holds $\sum_{j}r_{v,j}\times n_{{v,j}}=N$, since each vector space $H_{v}$ is isomorphic to $\mathbb{C}^{N}$. This in turn allows us to embed each $\mathrm{U}(n_{t(e),j})$ as a subgroup of the unitarities $\mathrm{U}(N)$ of $H_{t(e)}$ and motivates the following measure.
Definition 2.2.
Given an $N$-dimensional representation $R=\{(A_{v},H_{v}),(\phi_{e},U_{e})\}_{v\in Q_{0},e\in Q_{1}}$ of a connected quiver $Q$, we define the Dirac operator measure on $Q$ as
${\mathrm{d}}D:=\prod_{(v,w)\in Q_{0}\times Q_{0}}{\mathrm{d}}[D_{Q}(R)]_{v,w},\quad\text{ where }\quad{\mathrm{d}}[D_{Q}(R)]_{v,w}:=\prod_{e\in s^{-1}(v)\cap t^{-1}(w)}\prod_{j=1}^{l_{t(e)}}{\mathrm{d}}u_{e,j},$ (2.4)
being ${\mathrm{d}}u_{e,j}$ the Haar measure on $\mathrm{U}({{n_{t(e),j}}})$, where $u_{e,j}$ sits in the matrix $U_{e}$ associated to $e$ by $R$ in the respective block-diagonal entry in $U_{e}=\operatorname{\mathrm{diag}}(1_{r_{t(e),1}}\otimes u_{e,1},1_{r_{t(e),2}}\otimes u_{e,2},\cdots,1_{r_{t(e),l_{t(e)}}}\otimes u_{e,l_{t}(e)})$.
Therefore ${\mathrm{d}}D$ is the product Haar measure on $\prod_{e\in Q_{1}}\prod_{j=1}^{l_{t(e)}}\mathrm{U}({{n_{t(e),j}}})$ seen as subgroup of $\mathrm{U}(N)^{\#Q_{1}}$.
Definition 2.3.
The partition function of a given ${\mathscr{pS}}$-representation $R$ of a quiver $Q$ reads
$\mathcal{Z}_{Q,R}=\int\mathrm{e}^{-S(D)}{\mathrm{d}}D,\qquad N=\dim R\text{ and }D=D_{Q}(R).$ (2.5)
For any $\beta\in\Omega Q$, a Wilson loops [^5] We refer both to $\operatorname{\mathrm{hol}}\beta$ and to $\mathbb{E}[\operatorname{Tr}\operatorname{\mathrm{hol}}\beta]$ ambiguously as Wilson loops. is by definition
$\mathbb{E}[\operatorname{Tr}(\operatorname{\mathrm{hol}}\beta)]:=\frac{1}{\mathcal{Z}_{Q,R}}\int\operatorname{Tr}\operatorname{\mathrm{hol}}(\beta)\mathrm{e}^{-NS(D)}{\mathrm{d}}D.$ (2.6)
Remark 2.4.
For fixed $N$, the partition function $\mathcal{Z}_{Q}=\sum_{R\text{ ${\mathscr{pS}}$-rep of $Q$}}^{\dim R=N}\mathcal{Z}_{Q,R}$ seems to be a more interesting quantity, or even more so the sum over a class of quivers $Q$ encoding different background geometries, $\mathcal{Z}=\sum_{Q}\sum_{R\text{ ${\mathscr{pS}}$-rep of $Q$}}^{\dim R=N}\mathcal{Z}_{Q,R}$. For the moment we content ourselves with the partition function (2.5) for a fixed representation $R$ and a fixed quiver $Q$.
3. The Makeenko-Migdal loop equations for the spectral action
3.1. Notation
We now derive the constraints on the set of Wilson loops. We root an edge ${e_{\circ}}\in Q_{1}$ with $s({e_{\circ}})\neq t({e_{\circ}})$, and consider its associated unitarity $U_{e_{\circ}}$ as given by a fixed ${\mathscr{pS}}$-representation $R=\{(A_{v},H_{v}),(\phi_{e},U_{e})\}_{v,e}$ of $Q$. Now consider the infinitesimal variation of the spectral action by the change of variable exclusively for the unitarity $U_{e_{\circ}}$ at the edge ${e_{\circ}}$ as follows. Let
$U_{e_{\circ}}\mapsto U_{e_{\circ}}^{\prime}=\mathrm{e}^{\imath Y}U_{e_{\circ}},\qquad\imath Y\in\mathfrak{su}(N),\qquad\imath=\sqrt{-1},$ (3.1)
where $Y$ is given in terms of arbitrary matrices $\imath y_{k}\in\mathfrak{su}(n_{t(e),k})$ for $k=1,2,\ldots,l_{t({e_{\circ}})}=:L$ by
$\mathrm{e}^{\imath Y}:=\operatorname{\mathrm{diag}}\big{[}1_{r_{t(e),1}}\otimes\exp({\imath y_{1}}),1_{r_{t(e),2}}\otimes\exp({\imath y_{2}}),\ldots,1_{r_{t(e),{L}}}\otimes\exp(\imath y_{L})\big{]}.$
One should keep in mind that this implies also the substitution $U_{e_{\circ}}^{*}\mapsto(U_{e_{\circ}}^{\prime})^{*}=U_{e_{\circ}}^{*}\mathrm{e}^{-\imath Y}$, as it follows from the change (3.1). This rule defines a new representation $R^{\prime}$ differing from $R$ only by the value of the unitarities at the edge ${e_{\circ}}$, that is $R^{\prime}=\{(A_{v},H_{v}),(\phi_{e},\exp(\delta_{e,e_{0}}\imath Y)U_{e}\}_{v\in Q_{0},e\in Q_{1}}$, where $\delta_{e,e^{\prime}}$ is the indicator function on the edge-set.
Assume that along a given path $\gamma$ the combinations $e\bar{e}$ and $\bar{e}e$ are absent for each edge $e\in\gamma$. We call this type of paths reduced (Fig. 1) and it is trivial to see that reduction of a path (i.e. removing those pairs) yields a new one with unaltered holonomy. Consider then a reduced loop $\gamma$ that appears in the spectral action and contains the rooted edge ${e_{\circ}}$. This assumption allows (w.l.o.g. due to cyclic reordering) the decomposition
Figure 1. (a) An arbitrary (non-reduced, length-16) path on a rectangular lattice is shown. The ‘chair legs’ are edges present in the combination $e\bar{e}$. (b) Its reduced version. Notice that both paths have the same holonomy, though the path in (a) is larger than (b) in eight (the four removed combinations $e_{i}\bar{e}_{i}$ at the legs).
$\gamma=e_{\circ}^{\epsilon_{1}}\alpha_{1}e_{\circ}^{\epsilon_{2}}\alpha_{2}\cdots e_{\circ}^{\epsilon_{m}}\alpha_{m}$ (3.2a)
(cf. Fig. 2) where each of $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{n}\in\{+1,-1\}$ is a sign. This convention means that $e_{\circ}^{\epsilon}={{\bar{e}_{\circ}}}$ is the edge ${e_{\circ}}$ backwards if $\epsilon=-1$, while of course $e_{\circ}^{\epsilon}$ is ${e_{\circ}}$ itself if $\epsilon=1$. (The condition that $\gamma$ starts with ${e_{\circ}}$ implies $\epsilon_{1}=1$ above, but leaving this implicit is convenient.) By asking that each subpath $\alpha_{1},\ldots,\alpha_{m}\subset\gamma$ does not contain neither ${e_{\circ}}$ nor ${{\bar{e}_{\circ}}}$, one uniquely determines them. For another loop $\beta$, which also starts with ${e_{\circ}}$, under the same assumption that ${e_{\circ}}$ and ${{\bar{e}_{\circ}}}$ do not appear consecutively in $\beta$ in any order, a similar decomposition holds
$\beta$ $=e_{\circ}^{\sigma_{1}}\mu_{1}e_{\circ}^{\sigma_{2}}\mu_{2}\cdots e_{\circ}^{\sigma_{p}}\mu_{p}$ (3.2b)
in terms of signs $\sigma_{j}\in\{-1,+1\}$ and paths $\mu_{j}$ not containing neither the rooted edge ${e_{\circ}}$ nor ${{\bar{e}_{\circ}}}$. The only difference in notation —which we will keep throughout— is that $\gamma$ will refer to generalised plaquettes (i.e. contribution to the spectral action) while $\beta$ will be the path of a Wilson loop.
Figure 2. The most general reduced path $\gamma=e_{\circ}^{\epsilon_{1}}\alpha_{1}e_{\circ}^{\epsilon_{2}}\alpha_{2}\cdots e_{\circ}^{\epsilon_{m}}\alpha_{m}$ containing $e_{\circ}$ and/or ${{\bar{e}_{\circ}}}$ is shown here (omitting orientations for sake of simplicity). The most general Wilson loop can also be decomposed in subpaths in a similar way.
Take again the polynomial $f(x)=f_{0}+f_{1}x+f_{2}x^{2}+\ldots+f_{d}x^{d}$, and rephrase the spectral action of eq. (2.3) as
$S(D_{Q})=\operatorname{Tr}f(D_{Q})=\sum_{\begin{subarray}{c}\gamma\in\Omega Q\\ \gamma\text{ reduced }\end{subarray}}g_{\gamma}\operatorname{Tr}\operatorname{\mathrm{hol}}\gamma.\qquad$ (3.3)
Now $g_{\gamma}$ is a function of $f_{\ell(\gamma)}$ but possibly also of $f_{\ell(\gamma)+2},f_{\ell(\gamma)+4},\ldots$, whenever these last coefficients are non-zero. The reason for the contribution of the higher coefficients is the possible appearance of a contiguous pair of edges $e,\bar{e}$, for which the respective unitarities will satisfy $U_{e}U^{*}_{e}=1=U_{e}^{*}U_{e}$. These cancellations are not detected by the holonomy, which is the criterion used in (3.3) to collect all terms (instead of using, as in eq. (2.3), the $f_{0},\ldots,f_{d}$ coefficients and performing directly the sum over paths). For instance, if $\gamma$ is the path in Fig. 1 (b), then $g_{\gamma}$ depends on $f_{\ell(\gamma)}$ and $f_{\ell(\gamma)+8}$, since Fig. 1 (a) contributes the same to the spectral action. The function $g_{\gamma}=g_{\gamma}(f_{0},f_{1},\ldots,f_{d})$ is of course quiver-dependent.
3.2. Main statement
The Makeenko-Migdal or loop equations we are about to generalise appeared first in lattice quantum chromodynamics [MM79]. They have been a fundamental ingredient in the construction of Yang-Mills theory in two dimensions [Lév17, DGHK17, CPS23] in rigorous probabilistic terms.
Theorem 3.1 (Makeenko-Migdal equations for spectral action on quivers).
Let $R$ be a representation of a connected quiver $Q$ and let $N=\dim R$. Root an edge ${e_{\circ}}$ of $Q$ and abbreviate by $U=U_{{e_{\circ}}}$ the unitarity that $R$ determines for ${e_{\circ}}$. Then for any reduced loop $\beta$ the following relation among Wilson loops holds :
$\mathbb{E}\bigg{[}\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}=+1\end{subarray}}^{p}\frac{1}{N}\operatorname{Tr}(U^{\sigma_{1}}\operatorname{\mathrm{hol}}\mu_{1}\cdots U^{\sigma_{j-1}}\operatorname{\mathrm{hol}}\mu_{j-1})\frac{1}{N}\operatorname{Tr}(U^{\sigma_{j}}\operatorname{\mathrm{hol}}\mu_{j}\cdots U^{\sigma_{p}}\operatorname{\mathrm{hol}}\mu_{p})$
$\,\,-\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}=-1\end{subarray}}^{p}\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\mu_{1}U^{\sigma_{2}}\operatorname{\mathrm{hol}}\mu_{2}\cdots U^{\sigma_{j-1}}\operatorname{\mathrm{hol}}\mu_{j-1})\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\mu_{j}U^{\sigma_{j+1}}\cdots U^{\sigma_{p}}\operatorname{\mathrm{hol}}\mu_{p})\bigg{]}$
$=\sum_{\begin{subarray}{c}\gamma\in S(D)\\ \gamma\text{ \scriptsize reduced }\end{subarray}}g_{\gamma}\mathbb{E}\bigg{[}\sum_{\begin{subarray}{c}j=1\\ \epsilon_{j}=+1\end{subarray}}^{m}\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\beta\cdot U^{\epsilon_{j}}\operatorname{\mathrm{hol}}\alpha_{j}\cdots\operatorname{\mathrm{hol}}\alpha_{m}U^{\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}\cdots U^{\epsilon_{j-1}}\operatorname{\mathrm{hol}}\alpha_{j-1})$ (3.4)
$\qquad\qquad\qquad-\sum_{\begin{subarray}{c}j=1\\ \epsilon_{j}=-1\end{subarray}}^{m}\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\beta\cdot\operatorname{\mathrm{hol}}\alpha_{j}U^{\epsilon_{j+1}}\cdots\operatorname{\mathrm{hol}}\alpha_{m}U^{\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}\cdots\operatorname{\mathrm{hol}}\alpha_{j-1}U^{\epsilon_{j}})\bigg{]}.$
Notice that the second line takes the expectation value of $\tfrac{1}{N}\operatorname{Tr}(U^{\sigma_{1}}\operatorname{\mathrm{hol}}\mu_{1}\cdots U^{\sigma_{j-1}}\operatorname{\mathrm{hol}}\mu_{j-1}U^{\sigma_{j}})$$\times\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\mu_{j}\cdots U^{\sigma_{p}}\operatorname{\mathrm{hol}}\mu_{p})$, but $\sigma_{j}$ being $-1$ allows for a cancellation, hence the apparent lack of harmony between the two first lines of the lhs. We also stress that the first term in the lhs, which corresponds to $j=1=\sigma_{1}$, yields the input Wilson loop $\beta$ in the first trace and a constant path in the second; the latter yields a factor of $N$, which is cancelled by its prefactor.
Proof.
The loop or Dyson-Schwinger or Makeenko-Migdal equations follow from the identity
$\int\sum_{a,b=1}^{N}(\partial_{Y})_{a,b}\{(\operatorname{\mathrm{hol}}_{R^{\prime}})\beta_{b,a}\times\mathrm{e}^{-NS(D^{\prime})}\}{\mathrm{d}}D=0\qquad D^{\prime}=D_{Q}{(R^{\prime})}$ (3.5)
(The entries of the matrix derivative are $(\partial_{Y})_{a,b}=\partial/\partial Y_{b,a}$.) This identity follows from the invariance of the Haar measure at the rooted edge under the transformation (3.1), implying ${\mathrm{d}}D^{\prime}={\mathrm{d}}D$. Below, we show that this implies
$\mathbb{E}\bigg{[}\Big{(}\frac{1}{N}\operatorname{Tr}\otimes\frac{1}{N}\operatorname{Tr}\Big{)}(\partial_{Y}\operatorname{\mathrm{hol}}\beta)\bigg{]}=\mathbb{E}\bigg{[}\frac{1}{N}\operatorname{Tr}(\partial_{Y}S\operatorname{\mathrm{hol}}\beta)\bigg{]},$ (3.6)
and compute each quantity inside the trace(s). On the lhs, the matrix derivative is then the Rota-Stein-Turnbull noncommutative derivation
$\partial_{Y}Y^{k+1}=\sum_{l=0}^{k}Y^{l}\otimes Y^{k-l}\qquad(Y^{*}=Y\in M_{N}(\mathbb{C}),k=\mathbb{Z}_{\geq 0})$ (3.7)
while on the rhs, the matrix yields Voiculescu’s cyclic derivation, since the quantity it derives, $S$, contains a trace.
Recall that $\operatorname{\mathrm{hol}}$ is multiplicative, one has $\operatorname{\mathrm{hol}}\gamma=U^{\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}U^{\epsilon_{2}}\operatorname{\mathrm{hol}}\alpha_{2}U^{\epsilon_{3}}\cdots\operatorname{\mathrm{hol}}\alpha_{m-1}U^{\epsilon_{m}}\operatorname{\mathrm{hol}}\alpha_{m}$. With respect to the transformed representation $R^{\prime}$ we can compute the holonomy $\operatorname{\mathrm{hol}}_{R^{\prime}}\delta$ of any path $\delta$. This depends on $Y$ and ${e_{\circ}}$ but we use a prime in favor of a light notation and write only $\operatorname{\mathrm{hol}}^{\prime}\delta$. Since none of the subpaths $\alpha_{j}$ contains the transformed edges ${e_{\circ}}$ and ${{\bar{e}_{\circ}}}$, one has $\operatorname{\mathrm{hol}}^{\prime}\alpha_{j}=\operatorname{\mathrm{hol}}\alpha_{j}$, so
$\operatorname{\mathrm{hol}}^{\prime}(\gamma)=U^{{}^{\prime}\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}U^{{}^{\prime}\epsilon_{2}}\operatorname{\mathrm{hol}}\alpha_{2}U^{{}^{\prime}\epsilon_{3}}\cdots U^{{}^{\prime}\epsilon_{m}}\operatorname{\mathrm{hol}}\alpha_{m}$ (3.8)
where $U^{{}^{\prime}\epsilon}$ is $\mathrm{e}^{\imath Y}U$ if $\epsilon=1$ and $U^{*}\mathrm{e}^{-\imath Y}$ if $\epsilon=-1$. Therefore the variation of the loop $\gamma$ writes
$[\partial_{Y}\operatorname{Tr}\operatorname{\mathrm{hol}}^{\prime}\gamma]\big{|}_{Y=0}$ $=\imath\sum_{\begin{subarray}{c}j=1\\ \epsilon_{j}=+1\end{subarray}}^{m}U\operatorname{\mathrm{hol}}\alpha_{j}U^{\epsilon_{j+1}}\operatorname{\mathrm{hol}}\alpha_{j+1}\cdots U^{\epsilon_{n}}\operatorname{\mathrm{hol}}\alpha_{n}U^{\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}\cdots U^{\epsilon_{j-1}}\operatorname{\mathrm{hol}}\alpha_{j-1}$
$-\imath\sum_{\begin{subarray}{c}j=1\\ \epsilon_{j}=-1\end{subarray}}^{m}\operatorname{\mathrm{hol}}\alpha_{j}U^{\epsilon_{j+1}}\operatorname{\mathrm{hol}}\alpha_{j+1}\cdots U^{\epsilon_{n}}\operatorname{\mathrm{hol}}\alpha_{n}U^{\epsilon_{1}}\operatorname{\mathrm{hol}}\alpha_{1}\cdots U^{\epsilon_{j-1}}\operatorname{\mathrm{hol}}\alpha_{j-1}U^{*}.$
The cyclic wandering of any fix holonomy, say $\operatorname{\mathrm{hol}}\alpha_{1}$, in the rhs of the main result is due to Voiculescu’s cyclic derivation.
We now compute the variation of the Wilson line $\beta$, whose holonomy writes for the representation $R^{\prime}$ as $\operatorname{\mathrm{hol}}^{\prime}\beta=\prod_{j=1}^{p}(U^{\prime})^{\sigma_{j}}\operatorname{\mathrm{hol}}^{\prime}\mu_{j}=\prod_{j=1}^{p}(U^{\prime})^{\sigma_{j}}\operatorname{\mathrm{hol}}\mu_{j}$. To take the variation observe that $\operatorname{\mathrm{hol}}^{\prime}\beta$ is not inside a trace. For any matrices $A,B\in M_{N}(\mathbb{C})$, and $a,b,c,d=1,\ldots,N$,
$(\partial_{Y})_{b,a}[A\exp(\imath Y)B]_{c,d}\bigg{|}_{Y=0}$ $=\sum_{k=0}^{\infty}\frac{\imath^{k}}{k!}\sum_{l=0}^{k}A_{c,r}[Y^{l}\otimes Y^{k-1-l}]_{b,s|a,r}\bigg{|}_{Y=0}B_{s,d}=\imath A_{c,r}\delta_{b,s}\delta_{a,r}B_{s,d}.$
Using this rule for the previous expression of $\operatorname{\mathrm{hol}}^{\prime}\beta$, one obtains each summand for each occurrence of $U^{\pm 1}$ and the result follows after equating the indices $c=a$, and $b=d$, which is the initial situation in the initial identity (3.5). ∎
Suppose that the plaquettes in the action $S(D)=\sum_{\gamma}^{\text{\tiny reduced}}\tilde{g}_{\gamma}[\operatorname{Tr}\operatorname{\mathrm{hol}}\gamma+\operatorname{Tr}\operatorname{\mathrm{hol}}\bar{\gamma}]$, intersect each either ${e_{\circ}}$ and ${{\bar{e}_{\circ}}}$ exactly once. Notice that this time we have rewritten it as sum over pairs $\gamma$ and $\bar{\gamma}$ (which is always possible since the paths are in $\Gamma Q$ and the spectral action is real valued). Then
$\mathbb{E}\bigg{[}\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}=+1\end{subarray}}^{p}\frac{1}{N}\operatorname{Tr}(U\operatorname{\mathrm{hol}}\mu_{1}\cdots U^{\sigma_{j-1}}\operatorname{\mathrm{hol}}\mu_{j-1})\frac{1}{N}\operatorname{Tr}(U^{\sigma_{j}}\operatorname{\mathrm{hol}}\mu_{j}\cdots U^{\sigma_{p}}\operatorname{\mathrm{hol}}\mu_{p})$
$\,\,-\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}=-1\end{subarray}}^{p}\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\mu_{1}\cdots U^{\sigma_{j-1}}\operatorname{\mathrm{hol}}\mu_{j-1})\frac{1}{N}\operatorname{Tr}(\operatorname{\mathrm{hol}}\mu_{j}\cdots U^{\sigma_{p}}\operatorname{\mathrm{hol}}\mu_{p})\,\bigg{]}$
$=\sum_{\begin{subarray}{c}\gamma\in S(D)\\ \gamma=(U,\alpha)\text{ \scriptsize reduced }\end{subarray}}\tilde{g}_{\gamma}\mathbb{E}\big{[}\operatorname{Tr}(\operatorname{\mathrm{hol}}\beta\cdot U\operatorname{\mathrm{hol}}\alpha)-\operatorname{Tr}(\operatorname{\mathrm{hol}}\beta\operatorname{\mathrm{hol}}\bar{\alpha}\cdot U^{*})\big{]}.$ (3.9)
3.3. Graphical representation of the Makeenko-Migdal equations
We illustrate graphically the meaning of the Makeenko-Migdal equations. Let us place ${e_{\circ}}$ and the reversed edge ${{\bar{e}_{\circ}}}$ along a fixed axis of the picture. To represent a Wilson loop $\beta$ or a reduced generalised plaquette $\gamma$, we choose the following notation. In order to avoid drawings with several intersections, for each time that $\gamma$ or $\beta$ walks along either ${e_{\circ}}$ or ${{\bar{e}_{\circ}}}$, we jump to the next ‘plane’ in anti-clockwise direction around the fixed axis. Thus each of these planes represents abstractly the subpath $\mu_{j}\subset\beta$ or $\alpha_{j}\subset\gamma$ according to the decompositions (3.2a) and (3.2b), that is:
We kept a rectangular appearance for sake of visual simplicity, but the subpaths $\mu_{j}$ and $\alpha_{j}$ are arbitrary (as far as they have positive length). In fact, the depicted situation is still to some extent oversimplified, since the theorem describes the more general case that $\mu_{j}$ or $\alpha_{j}$ might be loops themselves (as $\alpha_{3}$ in Fig. 2), but this would render the pictures unreadable. The representation of the Makeenko-Migdal equations reads then as follows:
[…]
In the rhs, the very similar terms need a word of notation. The blue arrow is executed right after the green part of the path, while the red arrow follows only after the purple ones.
4. Applications and outlook
This last section has as aim to illustrate the power of the equations derived here when combined with the positivity conditions. This combination, sometimes known as ‘bootstrap’, appeared in [AK17] for lattice gauge theory and [Lin20] in a string context (for hermitian multimatrix models).
4.1. Positivity constraints
Let $v\in Q_{0}$ be fixed for this subsection and fix a representation $R$ of $Q$ of dimension $N$. Consider a complex variable $z_{\beta}$ for each loop $\beta$ based at a fixed vertex $v\in Q_{0}$, $z=\{z_{\beta}:\beta\in\Omega_{v}(Q)\}$, as well as the matrix
$P(z):=\sum_{\begin{subarray}{c}\beta\in\Omega_{v}(Q)\end{subarray}}z_{\beta}\operatorname{\mathrm{hol}}\beta,\qquad P(z)\in M_{N}(\mathbb{C}).$ (4.1)
It follows that $\operatorname{Tr}\big{[}P(z)P(z)^{*}\big{]}=\sum_{\beta,\alpha}z_{\beta}z^{*}_{\alpha}\operatorname{\mathrm{hol}}\beta\cdot(\operatorname{\mathrm{hol}}\alpha)^{*}=\sum_{\beta,\alpha\in\Omega_{v}(Q)}z_{\beta}z^{*}_{\alpha}\operatorname{\mathrm{hol}}(\beta\bar{\alpha})\geq 0$ independently of the $z$-tuple; this is preserved by expectation values, i.e.
$\sum_{\beta,\alpha\in\Omega_{v}(Q)}z_{\beta}z^{*}_{\alpha}\mathbb{E}[\operatorname{\mathrm{hol}}(\beta\bar{\alpha})]\geq 0,\qquad\text{for all }z\in\mathbb{C}^{\Omega_{v}(Q)},$ (4.2)
which is an equivalent way to state the positivity $\mathcal{M}\succeq 0$ of the matrix $\mathcal{M}\in\mathbb{C}[\![N,f_{0},f_{1},\ldots,f_{d}]\!]$ whose entries are given by
$(\mathcal{M})_{i,j}:=\mathbb{E}[\operatorname{\mathrm{hol}}(\beta_{i}\bar{\beta}_{j})]$ (4.3)
for any ordering of the loops $\{\beta_{1},\beta_{2},\ldots\}$ at the fixed vertex $v$. The positivity of $\mathcal{M}$ is clearly independent of the way we order these loops, as a conjugation by a permutation matrix (which is a unitary transformation) will not change the eigenvalues of $\mathcal{M}$.
Figure 3. The matrix entries for a $2\times 2$ submatrix of $\mathcal{M}$ (before taking expectation values) are shown in this figure. On the diagonal the trivial entries one sees result from the (holonomy of both) loops cancelling, which has the holonomy of the constant path. On the off-diagonal entries, one sees a non-trivial loop composition that goes first around the orange arrows and then along the green ones.
The choice for the matrix (4.3) is originally from [KZ24], who pushed forward the bootstrap for lattice Yang-Mills theory. The techniques of [Lin20] were implemented for fuzzy spectral triples for an interesting kind of hermitian matrix [HKP22] and a hermitian 2-matrix model [KP24]. The loop equations of [AK17] have been extended here to include arbitrary plaquettes that go along the rooted edge more than once, and Wilson loops that are allowed to do the same.
Example 4.1.
[…]
$S(D)=(f_{0}+2f_{2})N+x[\operatorname{Tr}\zeta+\operatorname{Tr}\zeta^{-1}]$ (4.4)
where we set $x=3f_{3}$. Since the terms in the even coefficients are just constants that when evaluating the Wilson loops expectation values, they disappear when dividing by the partition function and play no role in integration we set $f_{0}=f_{2}=0$. Now pick a loop $\beta=\zeta^{n}$ for positive $n\in\mathbb{Z}$. According to the loop equations (3.9), one has
$\mathbb{E}\bigg{[}\sum_{k=0}^{n-1}(\frac{1}{N}\operatorname{Tr}\otimes\frac{1}{N}\operatorname{Tr})(\zeta^{k}\otimes\zeta^{n-k})\bigg{]}=\frac{x}{N}(\mathbb{E}\operatorname{Tr}\zeta^{n+1}-\mathbb{E}\operatorname{Tr}\zeta^{n-1}).$ (4.5)
Defining the large-$N$ moments by $m_{j}:=\lim_{N\to\infty}\mathbb{E}[\tfrac{1}{N}\operatorname{Tr}\zeta^{j}]$ for each $j\in\mathbb{Z}$, this means
$\sum_{l=0}^{n-1}m_{l}\cdot m_{n-l}=x(m_{n+1}-m_{n-1}),\qquad(N\to\infty)$ (4.6)
since large-$N$ factorisation holds, $N^{-2}\mathbb{E}[\operatorname{Tr}\zeta^{i}\operatorname{Tr}\zeta^{j}]\to m_{i}\cdot m_{j}$, as $N\to\infty$. For the loop $\beta=\zeta^{-n}$ with $n\in\mathbb{Z}_{>0}$, one has
$-\sum_{j=0}^{n-1}m_{-(n-j)}\cdot m_{-j}=x(m_{-(n-1)}-m_{-(n+1)})\,,$ (4.7)
Finally, going through the derivation of the MM equations for the constant Wilson loop, one obtains the vanishing of the lhs, so $0=x(m_{1}-m_{-1})$, hence $\bar{m}_{1}=\overline{\mathbb{E}[\operatorname{Tr}\zeta]}={\mathbb{E}[\operatorname{Tr}\bar{\zeta}]}={\mathbb{E}[\operatorname{Tr}\zeta^{-1}]}=m_{-1}=m_{1}$, so $m_{1}$ is real (this can be derived by other means, but the loop equations yield this explicitly). Together with the last equation this implies $m_{-j}=m_{j}$ for all $j=1,2,\ldots$ and the moments can be arranged in the following (Toeplitz) matrix.
$\mathcal{M}=\begin{bmatrix}1&m_{1}&m_{2}&m_{3}&\ldots&\\ m_{-1}&1&m_{1}&m_{2}&\ldots&\\ m_{-2}&m_{-1}&1&m_{1}&\ldots&\\ m_{-3}&m_{-2}&m_{-1}&1&\ldots&\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}=\begin{bmatrix}1&m_{1}&m_{2}&m_{3}&\ldots&\\ m_{1}&1&m_{1}&m_{2}&\ldots&\\ m_{2}&m_{1}&1&m_{1}&\ldots&\\ m_{3}&m_{2}&m_{1}&1&\ldots&\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (4.8)
Thanks to Theorem 3.1, $\mathcal{M}$ can be computed recursively in terms of $y:=m_{1}$ and the coupling $x$,
$m_{1}$ $=y$ $m_{4}$ $=\frac{4y}{x}+\frac{3y^{2}}{x^{2}}+\frac{1}{x^{2}}+\frac{y}{x^{3}}+1$
$m_{2}$ $=\frac{y}{x}+1$ $m_{5}$ $=y+\frac{3y^{2}}{x}+\frac{2y^{3}}{x^{2}}+\frac{3}{x}+\frac{9y}{x^{2}}+\frac{6y^{2}}{x^{3}}+\frac{1}{x^{3}}+\frac{y}{x^{4}}$
$m_{3}$ $=y+\frac{y^{2}}{x}+\frac{1}{x}+\frac{y}{x^{2}}$ $m_{6}$ $=\frac{9y}{x}+\frac{18y^{2}}{x^{2}}+\frac{10y^{3}}{x^{3}}+\frac{6}{x^{2}}+\frac{16y}{x^{3}}+\frac{10y^{2}}{x^{4}}+\frac{1}{x^{4}}+\frac{y}{x^{5}}+1.$
Figure 4. Colored (in gray tones, if printed) are shown regions that violate $\mathcal{M}(x,y)\succeq 0$ as a consequence of the minors $\det(\mathcal{M}_{n}(x,y))<0$ from $n=1$ (tautological, not drawn), $n=2$ (the complement to the stripe $|y|<1$), $n=3$ (orange), $n=4$ (yellow), $n=5$ (green) and $n=6$ (blue). Only in the narrow white region satisfies the simultaneous conditions $\mathcal{M}_{n}(x,y)\succeq 0$, $n=1,2,\ldots,6$.
The positivity condition $\mathcal{M}(x,y)\succeq 0$ can be plotted on the first-moment–coupling plane in terms of the simultaneous positivity of the minors. The observed situation for a large class of hermitian matrix integrals are tight constraints for the first moment in terms of the coupling, which, by increasing the size of the minors, typically determine a curve $y=y(x)$—and by the respective loop equations, all the moments and thus the solution of the model. In this letter we do not claim convergence, but the results of this academic example encourage us to explore this combination in future works, including also a hermitian scalar (‘Higgs’) field that self-loops yield.
Acknowledgements
This work was mainly supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No818066) and also by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1-390900948 (the Heidelberg Structures Cluster of Excellence). I thank the Erwin Schrödinger International Institute for Mathematics and Physics (ESI) Vienna, where this article was finished, for optimal working conditions and hospitality. I acknowledge the kind answers by the group of Masoud Khalkhali at Western U., specially by Nathan Pagliaroli, on a question about bootstrapping.